please help me with this problem. Thanks sooooo much for your help!
Simplify. Write all the exponents in the denominator.
X^5y^2/x^3y^5

Found 2 solutions by saifoo.khan, stanbon:
Answer by saifoo.khan(162)

Hi,
x^2 / y^3
================
Cheers,
Saifoo.khan
einstien.saifoo@gmail.com

Answer by stanbon(75887)

Write all the exponents in the denominator.
X^5y^2/x^3y^5
-----
= (x^5/x^3)(y^2/y^5)
----
= (1/x^-2)(1/y^3)
------
= 1/(x^-2y^3)
=========================
Cheers,
Stan H.
=============